OpenStudy (nottim):

The base ethylamine, C2H5NH2 , has a pKb equal to 3.30. Calculate the hydroxide concentration, the pH, and the % ionization for a 0.42 M solution

OpenStudy (nottim):

I figured out Kb was 0.000501187, and I drew an incomplete ICE TAle. That's as far as I got.

OpenStudy (aaronq):

okay, thats good. have you written an equilibrium expression for the dissociation?

OpenStudy (nottim):

C2H5NH2--->HC2H5NH2+OH

OpenStudy (nottim):

ignored water, but the expression I think is [HC2H5NH]*[OH]/[C2H5NH2]

OpenStudy (aaronq):

don't forget Kb. Kb=[HC2H5NH]*[OH]/[C2H5NH2] so you said you had an ice table? you're gonna end up with: Kb=x^2/[C2H5NH2-x] so: 0.000501187=x^2/(0.42-x)

OpenStudy (nottim):

Why is it 0.42?

OpenStudy (aaronq):

it's given in the question, it's the concentration of ethyl amine before it abstracts a proton from water.

OpenStudy (nottim):

Oh gosh forgot to write that down in my work.

OpenStudy (aaronq):

thats probably why you were stuck :S. does it make sense now?

OpenStudy (nottim):

Just one more thing until I continue, can the 0.42 be ignored? Someone told me that before, but I don't know if they are right.

OpenStudy (aaronq):

the 0.42 cannot be ignored but the x in "0.42-x" could be. it depends on the "5% rule", because the Kb is so small, the amount ionized will be insignificant on 0.42 you can check this after by verifying that x is less than 5% of 0.42

OpenStudy (nottim):

Thanks!

OpenStudy (aaronq):

no problem, dude !